OpenStudy (anonymous):

There were 150 freshman and sophomores at the high school musical Reese and her friends attended. Freshman tickets were $2.00 each and sophomore tickets were $3.00 each. The receipts for freshman and sophomores tickets totaled $410. How many sophomores bought tickets? 40 90 110 120

OpenStudy (anonymous):

(150-x)*2+x*3 = 410 300 -2x+3x = 410 300+x = 410 x = 110